OpenStudy (tia_sweetheart17):

42xy/50yz times 10z/7y^2

razor99 (razor99):

i can help thats algebra right

OpenStudy (tia_sweetheart17):

yeah

OpenStudy (anonymous):

(52xy/50yz)*(10z/7y²)= (52*10*y*x*z)/(50*7*z*y³) = 520x/350y²

OpenStudy (anonymous):

@Luis Rivera where was the "x" gone ?

OpenStudy (anonymous):

It may be a typo, but the final ans should be 6x/5y^2